OpenStudy (mindblast3r):

The ratio of male to female employees at a local bank is 4:3. If 18 females work at this bank, how many males does the bank employ?

OpenStudy (mindblast3r):

@OrangeMaster

OpenStudy (anonymous):

24 i think

OpenStudy (mindblast3r):

Yes, that's the answer 100% but how did you get this?

OpenStudy (anonymous):

simple number of female employees=18=3/7 of the total no. of employees. let total no. of employees be X then 3/7(X)=18 =>X=42 So 4/7 of 42 is 24

OpenStudy (mindblast3r):

3/7?

OpenStudy (mindblast3r):

3 = 3 female employees, but what is the 7?

OpenStudy (anonymous):

ratio is 4:3 add 4&3 u get the denominator 7

OpenStudy (mindblast3r):

ahh.

OpenStudy (anonymous):

fell free to give a medal

OpenStudy (mindblast3r):

i will don't worry lol, right after you help me out so i can understand this completely :D

OpenStudy (mindblast3r):

"then 3/7(X)=18" what you mean by this?

OpenStudy (anonymous):

(7x18)/3=7x6=42

OpenStudy (anonymous):

so u get X=42

OpenStudy (anonymous):

good morning by the way

OpenStudy (mindblast3r):

lol you too, but i may live someone in like australia for all you know so it may be night here, but here it's morning time, so good morning lol

OpenStudy (mindblast3r):

"(7x18)/3=7x6=42" 7x18 = 126 / 3 = 42. But why did you put 7x6 now???

OpenStudy (anonymous):

because its easy to multiply 7&6 factors of 18 are 3x3x2 so you are basiclly doing 7x(3x2)x3/3=7x6=42

OpenStudy (anonymous):

its just a faster way of calculating

OpenStudy (mindblast3r):

umm..

OpenStudy (mindblast3r):

factors of 18?

OpenStudy (mindblast3r):

what does that mean?

OpenStudy (mindblast3r):

(really sorry lol) I can make this into 2 questions if you like so you can get more than 1 medal, cause your spending a lot of time helping me.

OpenStudy (anonymous):

factors mean the basic numbers it is made of

OpenStudy (mindblast3r):

Isn't there several other ways to get this answer?

OpenStudy (anonymous):

like factors of 32 are 2x2x2x2x2

OpenStudy (mindblast3r):

So it's like an exponent then?

OpenStudy (anonymous):

na medal is not the priority my goal is u should learn something out of this

OpenStudy (anonymous):

its somewhat like that

OpenStudy (mindblast3r):

Awe <3 thats so sweet haha well I am learning. Thank you very much :D

OpenStudy (mindblast3r):

"so you are basiclly doing 7x(3x2)x3/3=7x6=42"

OpenStudy (mindblast3r):

first the 7 represents the total number of people

OpenStudy (anonymous):

yup

OpenStudy (mindblast3r):

"(3x2)" represents what?

OpenStudy (mindblast3r):

the original ratio of the 4:3 male to female employees?

OpenStudy (anonymous):

6 my dear

OpenStudy (mindblast3r):

but where does the 3 come from? Is it the ratio of 4:3 male to female employees?

OpenStudy (anonymous):

do 24/18 u get 4/3

OpenStudy (anonymous):

which 3

OpenStudy (mindblast3r):

Now when you say "do 24/18 u get 4/3" do you mean divide 24 by 18?

OpenStudy (mindblast3r):

OHHHHH now i get it! lolol

OpenStudy (anonymous):

yes "/" means divide

OpenStudy (mindblast3r):

1.33

OpenStudy (mindblast3r):

1/3 = .33, 33% correct?

OpenStudy (anonymous):

4/3 is same as 4:3

OpenStudy (mindblast3r):

so 4/3 = 1.33, 1.33 = 133%

OpenStudy (mindblast3r):

Yeah, your just turning it into a ratio.

OpenStudy (mindblast3r):

" 7x(3x2)x3/3=7x6=42" can you explain this one more time please?

OpenStudy (mindblast3r):

and isn't their multiple ways to get the answer?

OpenStudy (mindblast3r):

I did it another way and I got the answer 24 but I'm not sure if it was just a coincidence. and I was wondering if you could help me out with that one.

OpenStudy (mindblast3r):

This is what I did, the ratio is 4:3, 4 males for every 3 female employees

OpenStudy (mindblast3r):

so 18 / 3 (3 female employees) =6

OpenStudy (mindblast3r):

then I did 6 x 4 (4 male employees) 24

OpenStudy (mindblast3r):

You there?

OpenStudy (mindblast3r):

That's how I got the answer 24 not sure if that's a correct way of getting it though.

OpenStudy (mindblast3r):

Also I just learned another way 4/3 = 1.33 x 18 = 24 (rounded)

OpenStudy (anonymous):

just wait i am sending a pic from my note book

OpenStudy (mindblast3r):

sure thing :)

OpenStudy (mindblast3r):

lol wow your amazing, going out of your way to help me, can't even thank you enough

OpenStudy (anonymous):

always happy to help

OpenStudy (mindblast3r):

what does "h" stand for on the page?

OpenStudy (mindblast3r):

or is that an 'n"?

OpenStudy (anonymous):

thats not "n" its a X

OpenStudy (mindblast3r):

oh lol sorry

OpenStudy (anonymous):

my teacher always points that out

OpenStudy (mindblast3r):

points what out? o.O

OpenStudy (anonymous):

THe X and the n

OpenStudy (mindblast3r):

lol, your teacher can't tell the difference either?

OpenStudy (anonymous):

i dont know why its just an X in the lower case cursive

OpenStudy (mindblast3r):

ah.. I'm not really familiar with cursive, that's why :P

OpenStudy (mindblast3r):

Okay I understand this thank you and goodbye lol.

OpenStudy (anonymous):

bye Hawk is always happy to help

OpenStudy (anonymous):

it was nice working with u